a 51 kg student is in a 8 kg sled at rest at the top of a hill. the student slides down the hill and reaches a speed of 116 km/h at the bottom. How high is the hill? express as m and round off to one decimal place


1
Expert's answer
2020-07-23T08:46:00-0400

Solution.

"m_1=51kg;"

"m_2=8kg;"

"\upsilon=116km/h=32.2m/s;"

"\dfrac{(m_1+m_2)\upsilon^2}{2}=(m_1+m_2)gh;"

"\dfrac{\upsilon^2}{2}=gh;"

"h=\dfrac{\upsilon^2}{2g};"

"h=\dfrac{(32.2m/s)^2}{2\sdot9.81m/s^2}=52.8m;"

Answer: "h=52.8m."